what solution set will satisfy if the system of equation 4x+y=8 a2x=3y+11 are manipulated using substitution?

Answer by ewatrrr(24785)


Hi
Solving system by substition method ('manipulated'...never heard it put quite that way :)
4x+y=8 Or y = (-4x-8)
2x=3y+11
2x =3(-4x-8) + 11
2x = -12x - 24 + 11
14x = -13
x = -13/14 and y = -4*-13/14 - 8 = 52/14 - 112/14 = -60/14 = -30/7
y = -30/7
CHECKING our Answer***
2x = 3y+11
-13/7 = -90/7 + 77/7 = -13/7